OpenStudy (anonymous):

A child is creating a pyramid with building blocks. The top three levels include 4 blocks, 10 blocks, and 16 blocks. Part 1: How many blocks would be needed for a pyramid 25 levels tall? Part 2: Use complete sentences to explain how a sum of an arithmetic series was applied.

OpenStudy (anonymous):

anyone~?

OpenStudy (raden):

given a1 = 4 and its difference d = 6. then use the formula : Sn = n/2 (2a1 + (n-1)d) just put n=25, and the values above into that formula

OpenStudy (anonymous):

2400? @RadEn

OpenStudy (raden):

Sn = n/2 (2a1 + (n-1)d) S25 = 25/2 (2*4 + (25-1)6) S25 = 25/2 * (8 + 144) S25 = 25/2 * 152 S25 = 1900

OpenStudy (anonymous):

ah.. makes more sense.. I appreciate that @RadEn

OpenStudy (raden):

okay

OpenStudy (anonymous):

any good explanation for part 2 ? @RadEn

OpenStudy (anonymous):

I've been doing these for like 5 hours I'm just tired lol